# Score-Based Metropolis-Hastings Algorithms

### v1

@@ 4.2 Entropy Regularization
Moreover, our algorithm involves an entropy regularization term alongside the proposed loss function defined in Section 4.1. To illustrate its effectiveness, we present the following result as motivation.
Proposition 2 (Estimated Acceptance Function).
Let $x,x^{\prime}\in\mathcal{X}$, and let $M\geq 1$. Then, the acceptance function defined as:
$a_{M}(x^{\prime},x)=\min\left\{\frac{1}{M},\frac{1}{M}\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}\right\}$
satisfies the detailed balance condition in (2).
Proof.
Proof is provided in the appendix. ∎
@@
It follows from Proposition 2 that there are infinitely many solutions to the detailed balance condition, and the acceptance probabilities can become infinitesimally small. While each of these acceptance functions is a mathematically valid solution to the detailed balance condition, in practice, sampling with these low-acceptance probability functions can lead to always rejecting the proposal, resulting in a lack of convergence within a reasonable timeframe.
@@
Therefore, if we solely optimize the loss functions in (4.1) and (4.1), we risk converging to minima that are legitimate solutions but impractical due to their very low acceptance probabilities. To address this issue, we propose adding an entropy regularization term to our loss function. The modified loss function is given by:
$\mathcal{L}(a)$ $=\mathbb{E}\bigg{[}\left\|\nabla\log a(X^{\prime},X)-\nabla\log a(X,X^{\prime})\right.$
@@
$\;+\nabla\log q(X^{\prime}\mid X)\|^{2}\bigg{]}+\lambda\,\mathbb{E}\left[H(\log a(X^{\prime},X))\right],$ (6)
where
@@
and $\lambda\geq 0$ is the weight of the regularization term. This regularization term encourages acceptance functions with higher entropy acceptance probabilities, thereby avoiding minima that result in very low acceptance probabilities.
@@
(a) Original samples.
(b) ULA.
(c) Score RW.
(d) Score MALA.
(e) Score pCN.
Figure 7: Comparison of different methods on the Swiss Roll dataset.
@@
4.3 Implementation
The final loss function used to train the acceptance network in our score MH framework consists of two components: the primary loss term, which enforces the reversibility condition, and an entropy regularization term. The primary loss is designed to minimize the difference between the gradients of the log acceptance probability and the gradients of the log posterior and proposal distributions.
@@ 4.3 Implementation
and where $x^{\prime(i)}=\alpha v^{\prime}+(1-\alpha)\tilde{x}$, with $\tilde{x}\sim p$ and $v^{\prime}\sim q(\cdot\mid\tilde{x})$, where $\alpha\in[0,1]$ controls how much $x^{\prime}$ deviates from the data distribution towards the proposal. In the first epochs of training, we begin with smaller values of $\alpha$ to favor data from the true distribution, and then gradually increase $\alpha$ to estimate the acceptance function around the proposal region. This approach is motivated by the fact that the score function may not be well-estimated outside of the data region, and starting with smaller $\alpha$ helps prevent biasing the training with poorly learned scores.
We present the pseudo-code for the proposed loss function in Algorithm 1 and the training procedure in Algorithm 2. We made several design choices that we observed to be beneficial for learning stable acceptance probabilities:
@@
  • Employing smooth activation functions, such as GeLU or Softplus.
  • Implementing annealed sampling from the proposal distribution, where instead of directly sampling from the target distribution, we start from a weighted combination of the observed data samples and the proposal samples.
@@ 5 Empirical Results
In this section, we present the empirical results on benchmark datasets including (Moons, Pinwheel, S-curve, and Swiss Roll), to demonstrate the validity of our approach. The results are presented in Figures 456, and 7.
@@
Results Analysis. The results in Table 1 illustrate the performance of four sampling methods: ULA, Score-based RW, Score-based MALA, and Score-based pCN, across four datasets using three evaluation metrics: Wasserstein-1 (W1), Wasserstein-2 (W2), and Maximum Mean Discrepancy (MMD).
Overall, the Score RW method consistently achieves lower W1 and W2 values compared to ULA across all datasets, highlighting its efficiency in sample-based inference. This trend is especially noticeable on the Moons and Pinwheel datasets, where the W2 metric significantly outperforms that of the ULA method. For more complex datasets like Swiss Roll, the score MH algorithms show noticeable improvements, with Score RW and Score MALA achieving significantly lower W1 and MMD scores when compared to the other methods.
@@
The results demonstrate that Score MALA is more robust to variations in step size compared to ULA. As seen in Figure 8, the W1, W2, and MMD metrics consistently show lower sensitivity to changes in step size for Score MALA, indicating its stability across different settings.
@@ 6 Discussion
The main limitation of the method is the additional computational cost required to train the acceptance network. Furthermore, more advanced architectures need to be proposed to extend this framework to high-dimensional data. Additionally, further theoretical analysis is needed to provide more appropriate priors on which acceptance function should be chosen.
@@ Acknowledgments
Vahid Tarokh’s work was supported in part by the Air Force Office of Scientific Research under award number FA9550-20-1-0397.
@@
In this appendix, we first discuss an alternative approximation of the acceptance function that relies solely on the score functions, derived using the Taylor expansion of the log densities. Next, we consider sampling with score-based MALA for heavy-tailed distributions. Third, we provide the proofs for the theoretical results. Finally, we present additional details of the experiments.
@@ Appendix C Proofs
Let $x,x^{\prime}\in\mathcal{X}$. Suppose that the gradients on both sides are equal, i.e.,
@@
Therefore we have that $C=0$, which concludes the proof. ∎
@@
which confirms that the detailed balance condition holds in this case. ∎

### v2

@@ 4 Algorithm
In addition to the original objective function introduced in Equation (3.3), our algorithm involves an entropy regularization term. To illustrate its effectiveness, we present the following result as its motivation.
Proposition 3.
Let $a:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ and let $M\geq 1$ be a scaling factor. Then,
$\mathcal{L}(a)=0\implies\mathcal{L}(\frac{a}{M})=0$
@@
It follows from Proposition 3 that there are infinitely many solutions to the detailed balance condition, including cases where the acceptance probabilities can become infinitesimally small. While each of these acceptance functions is a mathematically valid solution to the detailed balance condition, in practice, sampling with these low-acceptance probability functions can lead to always rejecting the proposal, resulting in a lack of convergence within a reasonable timeframe. Moreover, these infinitesimal values often cause training instabilities in the neural network that parameterizes the acceptance function.
@@
Therefore, if we solely optimize the loss function in (3.3), we risk converging to minima that are legitimate solutions but impractical due to their very low acceptance probabilities. To address this issue, we propose adding an entropy regularization term to our loss function. The modified loss function is given by:
$\mathcal{L}_{r}(a)$ $=\mathbb{E}\bigg{[}\left\|\nabla\log a(X^{\prime},X)-\nabla\log a(X,X^{\prime})\right.$
@@
$\;+\nabla\log q(X^{\prime}\mid X)\|^{2}\bigg{]}+\lambda\,\mathbb{E}\left[H(a(X^{\prime},X))\right],$ (10)
where:
@@
and $\lambda\geq 0$ is the weight of the regularization term. This regularization term encourages acceptance functions with higher entropy acceptance probabilities.
@@
Implementation.
Typically, only samples from the target distribution $p$ are observed, and the true $\nabla\log p$ is unknown. However, score parameterization through a neural network has proven successful in practice. We assume that we have an approximation of the score function previously estimated through a model $\tilde{s}\left(x\right)\approx\nabla\log p\left(x\right)$. Note that this is often the case in practice where an estimated score function $\tilde{s}$ is used in conjunction with ULA for sampling. The final loss function used to train the acceptance network in our score MH framework consists of two components: the primary loss term, which enforces the reversibility condition, and an entropy regularization term. The primary loss is designed to minimize the difference between the gradients of the log acceptance probability and the gradients of the log posterior and proposal distributions.
@@
and where $x^{\prime(i)}=\alpha v^{\prime}+(1-\alpha)\tilde{x}$, with $\tilde{x}\sim p$ and $v^{\prime}\sim q(\cdot\mid\tilde{x})$, where $\alpha\in[0,1]$ controls how much $x^{\prime}$ deviates from the data distribution towards the proposal. In the first epochs of training, we begin with smaller values of $\alpha$ to favor data from the true distribution, and then gradually increase $\alpha$ to estimate the acceptance function around the proposal region. This approach is motivated by the fact that the score function may not be well-estimated outside of the data region, and starting with smaller $\alpha$ helps prevent biasing the training with poorly learned scores. We present the pseudo-code for the proposed loss function in Algorithm 1 and the training procedure in Algorithm 2. We made several design choices that we observed to be beneficial for learning stable acceptance probabilities:
@@
  • Employing smooth activation functions, e.g., GeLU.
  • Clipping the gradients of the log acceptance function and the gradient of the log densities. This enables numerical stability and is also motivated by assumptions  2 and  3.
@@
(a) MALA, $\tau=0.5$
(b) MALA, $\tau=0.7$
(c) MALA, $\tau=0.9$
(d) MALA, $\tau=1.1$
(e) ULA, $\tau=0.5$
(f) ULA, $\tau=0.7$
(g) ULA, $\tau=0.9$
(h) ULA, $\tau=1.1$
Figure 6: MNIST: comparing annealed MALA (top) and annealed ULA (bottom) across $\tau$ values $0.5$, $0.7$, $0.9$, and $1.1$. Results show that MALA is more robust to step size. Both algorithms were run for $100$ denoising steps, with an additional $1000$ steps at the smallest noise level. $\tau$ is a fixed parameter that controls the adaptive step size.
@@ 5 Empirical Results
In this section, we present empirical results on various datasets (Moons, Pinwheel, S-curve, Swiss Roll, and MNIST), to demonstrate the validity of our approach. Further results on extreme value distributions are in Appendix B.
@@
Results Analysis. The results in Table 1 illustrate the performance of four sampling methods: ULA, Score-based RW, Score-based MALA, and Score-based pCN, across four datasets using three evaluation metrics: Wasserstein-1 (W1), Wasserstein-2 (W2), and Maximum Mean Discrepancy (MMD). Figures 34, and  5 illustrate the results for low-dimensional data. Overall, score-based methods achieve lower W1 and W2 values compared to ULA across all datasets. This trend is especially noticeable on the Moons and Pinwheel datasets, where the W2 metric significantly outperforms that of the ULA method. For more complex datasets like Swiss Roll, the score MH algorithms show noticeable improvements, with Score RW and Score MALA achieving significantly lower W1 and MMD scores when compared to the other methods.
@@
The results demonstrate that Score MALA is more robust to variations in step size compared to ULA as seen in Figures 6,  7 and 9.
@@
Figure 9: MNIST data generation comparing annealed ULA and annealed score-based MALA. $\tau$ is a fixed parameter that controls the adaptive step size.
Diffusion Models Application.
We evaluate the performance of the the proposed method with denoising score matching applied to the MNIST dataset using annealed ULA and annealing with score-based MALA. Annealing  (Song and Ermon, 2019; Song et al., 2021) involves a sequence of noise scales, progressively refining the samples from high noise to low noise, thus improving the stability and efficiency of score-based sampling. Further details on the annealing process and its implementation are provided in the appendix. Our results presented in Figures 6 and  9 indicate that the proposed method exhibits greater stability with respect to step size variations compared to annealed ULA, as measured by the sensitivity parameter $\tau$. We present further details in Appendix E.
@@ 6 Discussion
The main limitation of the method is the additional computational cost required to train the acceptance network. Furthermore, more advanced architectures need to be proposed to extend this framework to higher-resolution image data. Additionally, further theoretical analysis is needed to provide more appropriate priors on which acceptance function should be chosen.
@@ Acknowledgments
Ahmed Aloui and Vahid Tarokh’s work was supported in part by the Air Force Office of Scientific Research under award number FA9550-20-1-0397.
@@
In this appendix, we first discuss an alternative approximation of the acceptance function that relies solely on the score functions, derived using the Taylor expansion of the log densities. Next, we consider sampling with score-based MALA for heavy-tailed distributions. Third, we provide the proofs for the theoretical results. Finally, we present additional details of the experiments and further empirical results.
@@ Appendix C Proofs
We aim to prove that
$\mathcal{L}(a)=0\iff a\in\mathcal{A}.$
(Reverse Direction:) Suppose $a\in\mathcal{A}$, meaning that for every $x,x^{\prime}\in\mathcal{X}$, $a(x^{\prime},x)$ satisfies
$\frac{a(x^{\prime},x)}{a(x,x^{\prime})}=\frac{p(x^{\prime})q(x\mid x^{\prime})}{p(x)q(x^{\prime}\mid x)}.$
Taking the logarithm on both sides,
$\log a(x^{\prime},x)-\log a(x,x^{\prime})=\log p(x^{\prime})-\log p(x)+\log q(x\mid x^{\prime})-\log q(x^{\prime}\mid x).$
Differentiating both sides with respect to $x$ and $x^{\prime}$, we obtain
$\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})=\nabla\log p(x^{\prime})-\nabla\log p(x)+\nabla\log q(x\mid x^{\prime})-\nabla\log q(x^{\prime}\mid x).$
as we have that:
$\mathcal{L}(a)=\mathbb{E}_{x\sim p,x^{\prime}\sim q(\cdot|x)}\bigg{[}\left\|\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})-\nabla\log p(x^{\prime})+\nabla\log p(x)-\nabla\log q(x|x^{\prime})+\nabla\log q(x^{\prime}|x)\right\|^{2}\bigg{]}$
This implies $\mathcal{L}(a)=0$.
(Forward Direction:)
Suppose that $\mathcal{L}(a)=0$, i.e., as the expectation is taken over a positive variable we have that almost surely, for every $x,x^{\prime}\in\mathcal{X}$
@@
Therefore we have that $C=0$. Hence, $a\in\mathcal{A}$, which concludes the proof. ∎
@@
Let ^$\hat{\mathcal{L}}(a)$ be the empirical SBM loss computed over $N$ i.i.d. training samples:
^$\hat{\mathcal{L}}(a)=\frac{1}{N}\sum_{i=1}^{N}\left\|\nabla\log a(x_{i}^{\prime},x_{i})-\nabla\log a(x_{i},x_{i}^{\prime})-\nabla\log p(x_{i}^{\prime})+\nabla\log p(x_{i})-\nabla\log q(x_{i}\mid x_{i}^{\prime})+\nabla\log q(x_{i}^{\prime}\mid x_{i})\right\|^{2}.$
The true expected loss is
$\mathcal{L}(a)=\mathbb{E}_{x\sim p,x^{\prime}\sim q(\cdot|x)}\left[\left\|\nabla\log a(x^{\prime},x)-\nabla\log a(x,x^{\prime})-\nabla\log p(x^{\prime})+\nabla\log p(x)-\nabla\log q(x\mid x^{\prime})+\nabla\log q(x^{\prime}\mid x)\right\|^{2}\right].$
All we need to prove is that the loss function is bounded in order to apply the standard uniform convergence bound from statistical learning theory (Mohri, 2018).
Proving the boundedness of the loss function. By Assumption 1, $\mathcal{X}$ is compact, which implies that $\mathcal{X}\times\mathcal{X}$ is also compact. Furthermore, by Assumptions 2 and 3, the gradient terms $\nabla\log a$, $\nabla\log p$, and $\nabla\log q$ are continuous over a compact domain. By the Extreme Value Theorem, they attain their maximum absolute values, denoted by constants $C_{1},C_{2},$ and $C_{3}$, respectively:
$\sup_{x,x^{\prime}\in\mathcal{X}}\|\nabla\log a(x^{\prime},x)\|\leq C_{1},\quad\sup_{x,x^{\prime}\in\mathcal{X}}\|\nabla\log p(x)\|\leq C_{2},\quad\sup_{x,x^{\prime}\in\mathcal{X}}\|\nabla\log q(x^{\prime}\mid x)\|\leq C_{3}.$
Thus, the SBM loss is bounded as follows:
$\mathcal{L}(a)\leq\mathbb{E}_{x\sim p,x^{\prime}\sim q(\cdot|x)}\left[(2C_{1}+2C_{2}+2C_{3})^{2}\right]=C_{4},$
where $C_{4}=(2C_{1}+2C_{2}+2C_{3})^{2}$ is a finite constant.
Applying standard uniform convergence bounds (Mohri, 2018), for all $a\in\mathcal{F}$, with probability at least $1-\delta$, we obtain
$\sup_{a\in\mathcal{F}}\left|\mathcal{L}(a)-\hat{\mathcal{L}}(a)\right|\leq 2\mathcal{R}_{N}(\mathcal{F})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{N}}\right),$
where $\mathcal{R}_{N}(\mathcal{F})$ is the Rademacher complexity of the function class $\mathcal{F}$. ∎
Proof of Proposition 3.
Suppose that $\mathcal{L}(a)=0$, Let $a_{M}=\frac{a}{M}$, we prove in Proposition 1 that $\mathcal{L}(a)=0$ is equivalent to $a\in\mathcal{A}$, hence we will prove that $a_{M}\in\mathcal{A}$.
@@
which confirms that the detailed balance condition holds in this case.
Therefore, $a_{M}\in\mathcal{A}$. Hence, $\mathcal{L}(\frac{a}{M})=0$, which concludes the proof. ∎
@@ D.1 Dataset Descriptions
  • MNIST: MNIST consists of grayscale images of handwritten digits (0-9), each of size $28\times 28$. The dataset contains $60,000$ training samples and $10,000$ test samples. For our experiments, we normalize pixel values to $[-1,1]$ and convert the images into PyTorch tensors.
@@ D.2 Neural Network Architectures
MNIST Architectures.
For the MNIST Score Net, we employ a time-dependent UNet architecture with an encoder and decoder blocks and we use Softplus activations functions. For further details on the Score Network, please refer to the MNIST tutorial available at the following GitHub repository, as we use the same architecture https://github.com/yang-song/score_sde. For its acceptance network we use a Siamese like and that takes two inputs $(x,x^{\prime})$, embeds them via convolutional and residual layers, and processes them alongside learned time and step-size embeddings. We use GeLU activation functions. These branches do not share weights. The representation are then fed to a common branch. The acceptance network also encodes the time and step size parameters.
Annealed ULA and MALA for MNIST.
We evaluate annealed ULA and annealed MALA for sampling high-quality image reconstructions. We employ a denoisnig score matchign appraoch as presented in  (Song and Ermon, 2019; Song et al., 2021).
@@ D.3 Hyperparameters
Appendix E Additional Empirical Results
In this section, we include additional empirical results for three datasets. First, we perform a sanity check using a synthetic dataset, demonstrating that training the acceptance network with a learned score achieves similar performance to training with the true score, as shown in Figure 14. Furthermore, we present results for generating the Swiss Roll dataset in Figure 15. Finally, we provide additional results for the MNIST dataset, illustrating the robustness of Annealed MALA to the step size in Figure 16.
(a) Standard Random Walk Sampling.
(b) Score-based RW with true scores.
(c) Score-based RW with learned score.
(a) Original samples.
(b) ULA.
(c) Score RW.
(d) Score MALA.
(e) Score pCN.
Figure 15: Comparison of different methods on the Swiss Roll dataset.
Figure 16: Comparison of Wasserstein-1 (W1) values for Annealed MALA and Annealed ULA across different $\tau$ and step size settings. The plots classify regions based on whether the W1 values for both methods exceed or fall below a given threshold. Each subplot corresponds to a different threshold value ($30,32,34$). Colors indicate classification: (i) both methods above the threshold (blue), (ii) MALA below and ULA above (orange), (iii) ULA below and MALA above (green), and (iv) both methods below the threshold (gray). The x-axis represents the number of denoising steps, while the y-axis corresponds to the adaptive step size parameter $\tau$. We can see that annealed MALA exhibits a more robust behavior across different $\tau$ levels.